OpenStudy (anonymous):

A rifle is aimed horizontally at a target 50.0 m away. The bullet hits the target 2.90 cm below the aim point. Whats the bullet's flight time? Whats the bullet's speed as it left the barrel?

OpenStudy (anonymous):

You can see it like this: The time it takes the bullet to reach the target is how long it takes it to fall 2.9 cm due to gravity from rest:\[y(t) = 1/2 \times g \times t^{2}\] You solve that for t and get (assuming g=10 m/s^2) t = 0.076s flight time. For the speed you now know it flew 50m in 0.076s, and neglecting air friction that gives you 657.9m/s

OpenStudy (amistre64):

we could also time how long it take something to drop: .029 meters in free fall. h(t) = -4.9^2 + .029 = 0 .029 = 4.9 t^2 .029/4.9 = t^2 sqrt(.029/4.9) = t distance/time = speed; and lets assume the speed is constant. speed = 50(7)/sqrt(.29) = 350sqrt(.29)/.29 = 350sqrt(.29)/.29 =abt. 649.934; when we approximate with the exact time.

OpenStudy (amistre64):

the error of course is in how accurate you want to be :)